OpenStudy (anonymous):

Solve the equation. Show work. 7 = -V + 1

OpenStudy (osanseviero):

7 + V = -V + 1 +V 7 + V = 1 7 - 7 + V = 1 - 7 V = -6

OpenStudy (osanseviero):

Do you understand it?

OpenStudy (anonymous):

yes thank you! could you help me with a few more? they are all equations but they are different ya know

OpenStudy (osanseviero):

Ok, but I want to see you trying. It won't help you if I just solve it

OpenStudy (anonymous):

alright thank you. (Y + 7) + (-14) = 0 I know that first you have to get rid of the parenthesis i just dont know how to do that

OpenStudy (osanseviero):

(3 + 4) - (3) = (7) - (3) = 4 3 + (4 - 3) = 3 + (1) This means that the parenthesis that separate addition/substraction don't matter

OpenStudy (osanseviero):

So what should be your next step?